### v1

# The Hierarchical Growth of Bright Central Galaxies and Intracluster Light as Traced by the Magnitude Gap

Abstract
Using a sample of 2800 galaxy clusters identified in the Dark Energy Survey across the redshift range , we characterize the hierarchical assembly of Bright Central Galaxies (BCGs) and the surrounding intracluster light (ICL). To quantify hierarchical formation we use the stellar mass - halo mass (SMHM) relation for the BCG+ICL system and incorporate the magnitude gap (M14), the difference in brightness between the BCG (measured within 30 kpc) and 4th brightest cluster member galaxy within 0.5 . The inclusion of M14, which traces BCG hierarchical growth, increases the slope and decreases the intrinsic scatter in the SMHM relation, highlighting that it is a latent variable within the BCG+ICL SMHM relation. Moreover, the correlation with M14 decreases at large radii from the BCG’s centre. However, the stellar light within the BCG+ICL transition region (30 kpc - 80 kpc) most strongly correlates with the dark matter halo mass and has a statistically significant correlation with M14. As the light in the transition region and M14 are independent measurements, the transition region may grow as a result of the BCG’s hierarchical two-phase formation. Additionally, as M14 and ICL result from hierarchical growth, we use a stacked sample and find that clusters with large M14 values are characterized by larger ICL and BCG+ICL fractions, which illustrates that the merger processes that build the BCG stellar mass also grow the ICL. Furthermore, this may suggest that M14 combined with the ICL fraction can be used as a method to identify dynamically relaxed clusters.
1 Introduction
Bright Central Galaxies (BCGs) are massive, radially extended, elliptical galaxies located near the centre of their host cluster’s dark matter halo. BCGs grow hierarchically, by merging with smaller galaxies over time (e.g., De Lucia & Blaizot, 2007; Oser et al., 2010; van Dokkum et al., 2010). The combination of hierarchical formation and central location results in a correlation between the properties of the BCG and the cluster’s dark matter halo (Jones & Forman, 1984; Rhee & Latour, 1991; Lin & Mohr, 2004; Lauer et al., 2014). This galaxy-halo connection provides key information in characterizing the formation history of these massive galaxies.
Observations, simulations, and semi-analytic models (Croton et al., 2006; De Lucia & Blaizot, 2007; Guo et al., 2011; Tonini et al., 2012; Shankar et al., 2015) suggest that BCGs grow as a result of a two-phase formation (Oser et al., 2010; van Dokkum et al., 2010). At high redshift () a dense core (10kpc), which contains 25% of the stellar mass is formed via in-situ star formation. At , the outer envelope grows hierarchically as a result of major and minor mergers. This outer envelope extends to the faint and diffuse halo of intracluster light (ICL; Zwicky, 1933, 1951) that is observed around BCGs. Indeed, the presence of ICL can be used to differentiate BCGs from similarly massive non-central galaxies. Therefore, it is unsurprising that prior works have suggested that at least part of the ICL forms due to the BCG’s hierarchical assembly process (e.g., Murante et al., 2007; Golden-Marx et al., 2023).
Due to the faint and diffuse nature of the ICL, deep, high resolution photometry with a large field-of-view is needed to accurately characterize the ICL. Using such observations, individual clusters have measurements of the ICL that extend out to hundreds of kiloparsecs from the BCG (e.g., Kluge et al., 2020; Montes et al., 2021; Montes & Trujillo, 2022; Golden-Marx et al., 2023; Kluge et al., 2024). However, these observations can be enhanced by stacking the ICL signal surrounding many BCGs. Doing so, prior analyses have found that the ICL extends as far out as Mpc scales (Zibetti et al., 2005; Zhang et al., 2019c; Chen et al., 2022; Zhang et al., 2024). Although the ICL can be measured out to large radii, it is observationally challenging to distinguish the light associated with the BCG and that of the ICL. Therefore, we do not disentangle the BCG and ICL; instead, we focus on the BCG+ICL system and use fixed radial apertures to define our measurements, as introduced in Pillepich et al. (2018). In this analysis, we define similar radial regimes as Zhang et al. (2024); the BCG refers to the inner 30 kpc, the BCG+ICL transition regime is 30 kpc - 80 kpc, and the ICL is the light beyond 80 kpc.
The stellar mass-halo mass (SMHM) relation is a commonly used observational formalism to quantify the galaxy-halo connection for galaxy clusters ( / due to its low intrinsic scatter, , in stellar mass at fixed halo mass, which observations and simulations measure as 0.15 dex (e.g., Zu & Mandelbaum, 2015; Pillepich et al., 2018; Kravtsov et al., 2018; Golden-Marx & Miller, 2018). In logarithmic scale, this is a linear relation, which directly compares the stellar mass of the BCG to the total mass of the cluster with the form: log10(M∗) slope log10(Mhalo). Moreover, measurements of the SMHM relation’s parameters, such as the slope and provide insight into BCG growth and evolution (Gu et al., 2016; Golden-Marx & Miller, 2019; Golden-Marx et al., 2023).
As a result of the two-phase growth, information about the BCG’s recent stellar mass growth is contained within the BCG’s outer envelope and ICL (Oser et al., 2010; van Dokkum et al., 2010). It follows that measurements of the SMHM relation in both observations and semi-analytic models have found that the slope increases in progressively larger radii (Moster et al., 2018; Golden-Marx & Miller, 2019). Moreover, Golden-Marx et al. (2023) extended this analysis to the ICL and found that the SMHM relation’s slope increases to an asymptote at 100 kpc. Moreover, recent work found that the underlying halo mass is more strongly correlated with the light from the outer envelopes (50 kpc-100 kpc (Huang et al., 2022) or 30 kpc - 50 kpc (Golden-Marx et al., 2023)) than the light from the BCG’s core. Thus, the ICL is imprinted with information about recent mergers, which strengthens the correlation with the underlying halo mass.
Accounting for latent variables, which ideally increase the slope and decrease also strengthens statistical correlations. One observational measurement inherently tied to BCG hierarchical growth is the magnitude gap, the difference in band magnitude between the BCG and 4th brightest cluster member within half the radius enclosing 200 times the critical density of the Universe () (Dariush et al., 2010), referred to as M14 throughout this analysis. N-body simulations found that BCG stellar mass linearly increases with the number of progenitor galaxies (Solanes et al., 2016). Thus, as BCGs grow hierarchically, they increase in stellar mass and brightness, while the satellite galaxies remain fixed or merge with the BCG. Therefore, BCG growth increases M14 yielding the correlation between M14 and BCG stellar mass (Harrison et al., 2012; Golden-Marx & Miller, 2018, 2019; Golden-Marx et al., 2022). Thus, M14 traces the BCG’s hierarchical assembly (Golden-Marx & Miller, 2018) and acts as an indicator of the BCG’s dominance over other cluster member galaxies, similar to the Bautz-Morgan classification (Bautz & Morgan, 1970). Additionally, Vitorelli et al. (2018) found that clusters with larger magnitude gaps, defined using the 2nd brightest galaxy (M12), are more concentrated and likely formed earlier, suggesting that the magnitude gap traces cluster formation.
Based on Golden-Marx & Miller (2018), M14 is a statistical latent parameter within the cluster SMHM relation, as it increases the slope and decreases by as much as 50% (Golden-Marx & Miller, 2018, 2019; Golden-Marx et al., 2022). Moreover, reducing , tightens the constraints on the slope, which allowed for the detection of redshift evolution of the SMHM relation’s parameters (Golden-Marx & Miller, 2019; Golden-Marx et al., 2022). We also note that the M14 correlation holds if M12 is used (Golden-Marx & Miller, 2018). We use M14 because it yields a stronger correlation with early formation (Dariush et al., 2010). Moreover, the variance in the selection of the 4th brightest galaxy (and thus its magnitude) is less susceptible to large uncertainty due to foreground or background interloper galaxies, cluster mergers with multiple central galaxies, or recently infalling bright galaxies.
The ICL and M14 result from BCG hierarchical assembly. However, it remains unclear how these parameters are related. The ICL is a direct measurement of the diffuse light that surrounds the BCG and can be used as a fossil tracer of the BCGs assembly history, while M14 is a measure of the BCGs hierarchical assembly within the cluster environment. Therefore, in this paper we present measurements of the correlation of the BCG+ICL stellar content with halo mass and M14 (an extension of the work presented by Golden-Marx et al. (2023)) to determine for the first time how these parameters correlate in the context of BCG hierarchical assembly and the SMHM relation. For this analysis, we use the Dark Energy Survey (DES; DES Collaboration, 2005) Y6 sample of clusters, currently the only statistically large (> 1000) sample of clusters that such measurements of the ICL can be taken within the redshift range .
The remainder of this paper is outlined as follows. In Section 2, we define our observational sample, the DES Y3 redMaPPer cluster sample and DES Y6 data, and the data reduction methods. In Section 3, we present the measurements of the stellar mass content of the BCG+ICL system, halo mass, and M14. In Section 4, we describe the hierarchical Bayesian model used to measure the parameters of the SMHM relation. In Section 5, we present our analysis of the SMHM relation for the BCG+ICL system. In Section 6, we discuss the correlation between M14 and ICL, including the ICL fraction. Finally, we conclude in Section 7.
In this analysis, we assume a flat CDM universe, with =0.30, =0.70, H0=100 km/s/Mpc with =0.7.
2 data
2.1 DES Y6 data
The Dark Energy Survey (DES) is a wide-area multi-band photometric sky survey covering 5,000 square degrees across five photometric wavebands () taken using the Victor M. Blanco 4-m telescope beginning with Science Verification in 2012 and ending in 2019 with the Year 6 observations. The DES Year 6 data, which includes all prior DES data releases, features several layers of single-epoch images combined to create deep photometry that allow us to capture low surface brightness features. DECam, the imager used for DES, features a large 2.2 degree field-of-view, a 570 megapixel camera constructed using a CCD mosaic, and a low-noise electronic readout system (Flaugher et al., 2015). These properties make DECAM ideal for studying faint and extended diffuse sources, including the ICL, in the optical from the ground in a single field-of-view.
The data used throughout this analysis, both the source catalogs and the images, come from the DES Y6 co-addition (this includes publicly released data from DES Data Release 2 (Abbott et al., 2021)). The DESY6 images are deeper than the Y3 versions and provide a more uniform co-add with an improved background subtraction. Since this work focuses on the diffuse ICL, these improvements, which include the improved masking of faint galaxies located spatially near/within the ICL, are key. Such faint galaxies can unintentionally be treated as part of the ICL, thus, these improvements enhance our ability to accurately measure the contribution from the ICL by strengthening our ability to delineate between the diffuse light and total light within the cluster, which includes all cluster member galaxies and the ICL.
The improvements seen in the final images presented in Abbott et al. (2021) take advantage of additional exposures and tilings that cover the entire DES footprint. For these images, a single-epoch background subtraction is performed, instead of a “global” Swarp (Bertin, 2010) background subtraction. Like in Golden-Marx et al. (2023) and Zhang et al. (2024) the science images are an average of the photometry. Moreover, the DESY6 photometry allows for a fainter detection threshold (5 than what was used in previous data releases (10).
2.2 redMaPPer Cluster Sample
In this analysis, we use the DES Y3 redMaPPer cluster catalog version 6.5.22+2, based on DES Year 3 Gold photometric data (Sevilla-Noarbe et al., 2021). This version of the redMaPPer cluster catalogue contains over 21,000 galaxy clusters with a redMaPPer defined richness, , greater than 20, which corresponds to a halo mass lower limit of M⊙ (McClintock et al., 2019). Moreover, we take advantage of redMaPPer’s membership catalogue which provides the probability () and DES r-band model magnitude of each cluster member, to measure M14 and central galaxy probability (), for BCG identification. We note that redMaPPer selects a luminous cluster galaxy located nearest to the centre of the cluster’s gravitational potential well and provides the information for the five most probable central candidates per cluster. We assume that the most likely central galaxy candidate is the BCG. Based on multi-wavelength measurements, this selection has been shown to be correct with an frequency (Saro et al., 2015; Zhang et al., 2019a; Bleem et al., 2020).
2.3 Measuring the ICL
The BCG and ICL profiles used in this analysis are data products derived as part of Zhang et al. (2024), which uses a similar methodology as Zhang et al. (2019c), Sampaio-Santos et al. (2021), Golden-Marx et al. (2022), and Golden-Marx et al. (2023). Below, we briefly summarize the methodology used in Zhang et al. (2024).
We analyse only clusters in the redshift range 0.2 0.6, which reduces the number of redMaPPer ( 20) clusters from 21092 to 15654.
The BCGs were selected as the most probable central galaxy provided by redMaPPer. The impact of using all BCGs regardless of redMaPPer is discussed in the Appendix.
Using these BCGs, the DES database was queried to select images within a 0.15 deg 0.15 deg region centred on the BCG. These individual images were combined (using mean pixel values) to create the final coadd image centred on the BCG. The images used throughout this analysis have not had a local sky background subtraction (measured either with Source Extractor (Bertin & Arnouts, 1996) or Swarp (Bertin, 2010)) applied. Instead, these images underwent a global background subtraction process (estimated across the whole Field-of-View, about 2.5 for single exposure images; Bernstein et al., 2017). This background subtraction is done because local background regions often include the ICL, which results in an underestimation of the light in the BCG+ICL outer profiles (e.g., von der Linden et al., 2007; Golden-Marx & Miller, 2018).
To prevent contamination of our BCG+ICL light profiles from bright foreground stars and galaxies, clusters with BCGs near such objects were removed, reducing the number of available clusters in our sample to 7042.
Using the BCG as the centre, all objects covering a 0.2 deg 0.2 deg region surrounding the BCG in the DES database were identified. We acknowledge that miscentring exists in redMaPPer (see Zhang et al. (2019a)). However, this is not corrected for in the reduction process. As shown in the Appendix, when we look at high and low probability centrals, we find no evidence that it impacts the parameters of the SMHM relation as shown in Tables 4 and 5.
Elliptical masks, with a semi-major axis of 3.5 (Kron, 1980), were placed around all galaxies above a selected masking limit, excluding the BCG, allowing for the measurement of a BCG+ICL profile. The masking limit is a -band magnitude limit, based on the cluster’s photometric redshift, of 0.2, where is the characteristic luminosity of a cluster red galaxy luminosity function measurement (Zhang et al., 2019b).
Using the BCG as the central point, the surface brightness profiles (mean values) of the unmasked regions for each galaxy cluster, representative of the BCG+ICL light profile, were measured in circular radial annuli. We note that using simulations, Brough et al. (2024) find no statistical difference between measurements of the ICL done using circular and elliptical annular profiles. These measurements are the surface brightness radial profile for each cluster, which contain light from the cluster’s BCG and ICL as well as some background light. Masked regions are excluded from the measurement process.
The surface brightness measurements described in the previous step also contain residual background light and contributions from unmasked foreground and background objects. For each radial surface brightness measurement, we take their values at the radial range beyond 500 kpc as the “background” value, and subtract this “background” value from the corresponding measurements. The robustness of this background measurement was discussed in the Appendix of Golden-Marx et al. (2023).
Lastly, we remind the reader that unlike Zhang et al. (2019c), Sampaio-Santos et al. (2021), and Zhang et al. (2024), we do not stack the measurements of the ICL on a secondary cluster property such as richness. Instead, we only combine images of the same cluster to create a deeper photometric image.
3 The Observable Measurements for Constructing the SMHM Relation
3.1 BCG Stellar Mass
The stellar mass is estimated in the same manner as in Golden-Marx et al. (2023), so we summarize the method. For each cluster, using the BCG+ICL system’s light profile (see Section 2.3) we measure the apparent magnitude within a given radial aperture by integrating the light profile between the selected inner and outer radii in the DES i-band. In this analysis we use the following apertures: 0-30 kpc, 0-80 kpc, 0-150 kpc, 0-225 kpc, 0-300 kpc, 30-80 kpc, 80-150 kpc, 150-225 kpc, 225-300 kpc.
Using the selected aperture magnitudes, we use the EzGal (Mancone & Gonzalez, 2012) SED modelling software to estimate the stellar mass. For EzGal, we assume a passive spectral model, as DES photometry prevents us from statistically constraining additional star formation parameters such as burst times and formation epochs. For converting magnitude to stellar mass, we assume the same parameters as Golden-Marx & Miller (2019), Golden-Marx et al. (2022), and Golden-Marx et al. (2023); a Bruzual & Charlot (2003) stellar population synthesis model, a Salpeter (1955) Initial Mass Funtion (IMF), a formation redshift of , and a metallicity of 0.008 (66% Z⊙). These parameters were selected to minimize the statistic between the measured and EzGal modeled photometry. Following Golden-Marx et al. (2022) and Golden-Marx et al. (2023), we assume a subsolar metallicity for each BCG+ICL system since observations find that the metallicity of early type galaxies decreases radially outward (McDermid et al., 2015) and light from the ICL has been characterized as having a subsolar metallicity (Montes & Trujillo, 2018). Additionally, we note that our stellar mass estimate is independent of our choice of formation redshift and that changing the Salpeter (1955) IMF to a Chabrier (2003) IMF uniformly lowers our stellar mass estimates by 0.25 dex. Thus, the SMHM relation’s amplitude is the only SMHM parameter impacted.
Since our Bayesian infrastructure relies on an estimate of measurement uncertainty, we assume the same uncertainty in stellar mass, 0.06 dex, as in Golden-Marx et al. (2022) and Golden-Marx et al. (2023). This is likely a lower limit on the estimated uncertainty in stellar mass, in particular when we incorporate the ICL. If a larger uncertainty is required, this only changes . The remaining parameters are consistent. While our MCMC analysis measures , a detected change of with radius may reflect our underestimation of the uncertainty in the stellar mass at large radii.
3.2 DES Cluster Richnesses and Halo Masses
Unlike in Golden-Marx et al. (2023), which used the DES-ACT overlap sample (Hilton et al., 2021) and SZ-estimated halo masses, for the DES clusters, we use redMaPPer richness, . This richness is converted to halo mass using the DES Year 1 calibrated mass-richness relation (McClintock et al., 2019), given by Equation 1,
where refers to , is the redMaPPer photometric redshift and is the redMaPPer richness. We note that the McClintock et al. (2019) mass-richness relations has an intrinsic scatter associated with the halo mass at fixed richness. While not shown in Equation 1, we account for this scatter in our Bayesian MCMC analysis, as discussed in Section 4. Although we are using the DES Y3 redMaPPer catalogue a preliminary analysis of the DES-redMaPPer Y3 mass-richness relationship is consistent with the Y1 analysis.
We note that a bias has been found in the weak lensing calibration of the McClintock et al. (2019) mass-richness relation. The halo masses are biased 0.1 dex low, with the 30 clusters most strongly impacted (Abbott et al., 2020; Costanzi et al., 2021). For this analysis, if only the low- clusters have halo masses that are biased low, this bias would result in a value of the slope that is also biased low and may also impact the offset parameter. Since this would be a uniform bias across each aperture measurement of the SMHM relation, the bias would not affect our ability to characterize how incorporating ICL and M14 change the SMHM relation and thus would not impact the conclusions of this analysis.
Here, we treat the uncertainty in the mass-richness relation as a fixed value identical to what was found for the SDSS and DES redMaPPer clusters in Golden-Marx & Miller (2019) and Golden-Marx et al. (2022) because these measurements come from the same algorithm. We remind the reader that our chosen value, 0.087 dex, was determined in Golden-Marx & Miller (2019) from a joint analysis where the parameters for the SMHM relation were simultaneously determined for a sample with halo masses estimated by richness and the caustic phase-space technique. This scatter in halo mass at fixed richness corresponds to 0.20 in a natural log scale, which is in excellent agreement with Rozo et al. (2015), which measures a value between 0.17-0.21.
3.3 Magnitude Gap (M14)
M14 is defined uniformly throughout this paper, regardless of our choice of BCG+ICL aperture stellar mass. For the DES-redMaPPer clusters we identify cluster members as those with . This is a less restrictive than Golden-Marx & Miller (2019) and Golden-Marx et al. (2022) since we do not match to the C4 catalog (Miller et al., 2005). This choice optimizes the number of clusters in our sample and only reduces the sample from 7042 to 6903 clusters. When we did this analysis with a threshold of our results were consistent.
We define M14 following Dariush et al. (2010) as the difference in the band apparent model magnitude of the 4th brightest cluster member (with 0.9) within 0.5 and the BCG’s inner 30 kpc band apparent magnitude. We estimate for redMaPPer clusters using the following equations from Rykoff et al. (2014):
where is the redMaPPer richness, and is the redMaPPer cutoff radius, given by Equation 3:
Since we use DES data, we estimate the M14 uncertainty in the same manner as in Golden-Marx et al. (2022) and assume a value of 0.31. While this value may be an upper limit due to our lower criterion (as opposed to 0.984 used in Golden-Marx et al. (2022)) this is unlikely to change our results as our Bayesian analysis is robust to moderate changes in the M14 uncertainty. For example, if we increase the M14 uncertainty to 0.5, our values remain in agreement.
4 The Statistical Modelling of the SMHM Relation
We measure the parameters of the SMHM relation using a similar hierarchical Bayesian MCMC approach to what is described in Golden-Marx & Miller (2019) and Golden-Marx et al. (2023). We quantify the SMHM relation using Equation (4), when M14 is incorporated
and Equation 5, when it is not
where is the mathematical offset, is the slope of the SMHM relation, is the M14 parameter, and is the intrinsic scatter in stellar mass at fixed . This is a nested model, if =0, Equation 4 reverts to Equation 5. The only differences between the model used in this analysis and those in Golden-Marx & Miller (2019) and Golden-Marx et al. (2023) are that we do not account for redshift evolution and we include , respectively.
A Bayesian formalism works by convolving prior information associated with a selected model with the likelihood of the observations given that model, to yield the posterior distribution for the non-nuisance parameters in our model. To determine the posterior distributions for each SMHM parameter, our MCMC model generates values for the observed aperture stellar masses, richness-estimated halo masses, and M14s at each step in our likelihood analysis, which are directly compared to the observed measurements.
We model the log10 BCG stellar masses (), log10 halo masses (), and M14 values () as normal distributions with mean values (locations) taken from our measurements. The standard deviations associated with each value is taken from measurement uncertainties, described in Section 3 (, , ) as well as a stochastic component from a beta function, (Golden-Marx & Miller, 2018), which allows for additional uncertainty on the observational errors. These errors are treated statistically in the Bayesian model as free nuisance parameters , , and . Additionally, as done in Golden-Marx & Miller (2019), Golden-Marx et al. (2022), and Golden-Marx et al. (2023), we reduce the covariance between and by subtracting off the median values of the upper and lower limits in log/ and log/, 11.3 and 14.55 respectively. For consistency we subtract off the median stellar mass measured within 0-30 kpc in each aperture bin.
Due to the lack of detected evolution in the stellar mass of the ICL found in Golden-Marx et al. (2023) and Zhang et al. (2024), and the results from Golden-Marx et al. (2022), which find that the majority of the evolution in the parameters of the SMHM relation (when including M14) occurs in the redshift range , much of which is below the redshift range we observe, we do not allow for the parameters of the SMHM relation to vary with redshift.
The parameters of the SMHM relation are given in Table 1, where each cluster is a component in the summed log likelihood and the terms marked are representative of the observed data.
| Symbol | Description | Prior |
| The offset of the SMHM relation | (-20,20) | |
| The high-mass power law slope | Linear Regression Prior | |
| The stretch parameter which describes the stellar mass - M14 stratification | Linear Regression Prior | |
| The uncertainty in the intrinsic stellar mass at fixed | ||
| The underlying distribution in stellar mass | Deterministic (inferred from Equation 4) | |
| The underlying distribution | (14.11,) | |
| The underlying M14 distribution | (1.74,) | |
| The uncertainty between the observed stellar mass and intrinsic stellar mass distribution | 0.06 dex | |
| The uncertainty associated with log10() | 0.087 dex | |
| The uncertainty between the underlying and observed M14 distribution | 0.31 dex |
Additionally, we reiterate that this is a nested Bayesian model, which allows for to equal 0. Thus any statistically significant non-zero is real and signifies that including M14 improves the fit to the measured data. If a correlation between M14 and stellar mass did not exist in our data or added additional noise, we would not detect a statistically significant parameter. This discussion is expanded in Table 6 in the Appendix.
We express the entire posterior as:
5 The Observed SMHM Relation
As discussed in Section 1, many recent analyses have focused on the connection between the ICL and the cluster’s dark matter halo. Recent observations (e.g., Montes & Trujillo, 2019) suggest the light distribution of the ICL may trace the halo’s dark matter distribution. Moreover, prior results have identified statistically significant correlations between the stellar mass contained within the ICL and the cluster’s halo mass (e.g., Sampaio-Santos et al., 2021; Huang et al., 2022; Golden-Marx et al., 2023). As presented in Golden-Marx et al. (2023), accounting for the ICL increases the slope, , of the SMHM relation by 0.1-0.15. Thus, as a result of BCG+ICL hierarchical growth, the larger slope is measured because ICL, due to recent growth, is more correlated with the dark matter halo mass. Here, for the first time, we analyse the correlation between M14 and the ICL and investigate both quantitatively and qualitatively the impact of incorporating the ICL and M14 into the SMHM relation.
Our observational sample consists of all redMaPPer clusters ( > 20) in the redshift range with at least 4 high probability member galaxies ( > 0.90). We remove a large fraction of clusters that as a result of masking of either bright stars or nearby galaxies, we are unable to measure the BCG+ICL profile, reducing our sample size to 2929. Lastly, we remove those clusters that we were unable to measure the light out to 300 kpc leaving our final analysis sample as 2788 cluster. These criteria are summarized in Table 2.
| Criteria | Number of Clusters |
| > 20 redMaPPer sample | 21092 |
| 0.2 0.6 | 15654 |
| No nearby bright foreground objects | 7042 |
| 4 0.9 members | 6903 |
| ICL unimpacated by masking of nearby galaxies | 2929 |
| Can measure ICL profile out to 300 kpc | 2788 |
5.1 The BCG+ICL+M14 SMHM relation
First, we address the entire BCG+ICL profile and how the SMHM relation varies as we increase the outer radius. Specifically, we use the 0-30 kpc, 0-80 kpc, 0-150 kpc, 0-225 kpc, and 0-300 kpc apertures as shown in Figure 1. We note that the ICL is defined as incorporating all light beyond the transition region ( 80 kpc). However, as we were unsure of whether M14 continues to correlate with BCG+ICL stellar mass out to the full extent of the ICL, we divide the ICL into multiple apertures.
Figure 1 visualizes the stellar mass within the selected aperture plotted against the halo mass for each annulus, colour coded by M14. In each subplot, we overplot three lines, which are representative of the median of one-third of the data sorted by M14. The median M14 values are 0.90 (dark blue), 1.50 (green) and 2.10 (yellow). In the final subplot, we overlay the posterior distribution from our MCMC analysis for samples representative of the upper and lower M14 values to illustrate our fitting.
| Data | Inner Radius | Outer Radius | ||||
| Core | 0 | 30 | – | |||
| Core | 0 | 80 | – | |||
| Core | 0 | 150 | – | |||
| Core | 0 | 225 | – | |||
| Core | 0 | 300 | – | |||
| Core | 0 | 30 | ||||
| Core | 0 | 80 | ||||
| Core | 0 | 150 | ||||
| Core | 0 | 225 | ||||
| Core | 0 | 300 | ||||
| No Core | 0 | 30 | – | |||
| No Core | 30 | 80 | – | |||
| No Core | 80 | 150 | – | |||
| No Core | 150 | 225 | – | |||
| No Core | 225 | 300 | – | |||
| No Core | 0 | 30 | ||||
| No Core | 30 | 80 | ||||
| No Core | 80 | 150 | ||||
| No Core | 150 | 225 | ||||
| No Core | 225 | 300 |
To quantify the impact of including the ICL and M14 on the parameters of the SMHM relation, we present the values of the posterior distribution (both for when M14 is and is not incorporated) in Table 3. For consistency, we see that the results in the radial range of kpc are in excellent agreement with the results presented in Golden-Marx & Miller (2019) and Golden-Marx et al. (2022). Moreover, these SMHM relation parameters are within 1 of the values of and measured in Golden-Marx et al. (2023) (due to the different median subtraction values, we can’t compare ), which shows our results are not biased by the use of richness-inferred halo masses as opposed to SZ-estimated halo masses.
In terms of including M14, Figure 2 directly compares how and vary with outer radius when M14 is (blue) and is not (red) incorporated.
Across each aperture, we see clear trends such that the incorporation of M14 increases by 0.10-0.15 matching prior results (Golden-Marx & Miller, 2018, 2019; Golden-Marx et al., 2022). Moreover, reaches an asymptote at a radii of 150 kpc, in strong agreement with the trends found in Golden-Marx & Miller (2019) and Golden-Marx et al. (2023). Additionally, the inclusion of the ICL increases by 0.1-0.15 as found in Golden-Marx et al. (2023) (a similar increase to what is found by including M14). As in Golden-Marx & Miller (2019), the value asymptotes at is significantly higher when M14 is accounted for, by approximately 0.1.
The change of is more nuanced. When M14 is incorporated we find that decreases, in agreement with the trends for the BCG (Golden-Marx & Miller, 2018, 2019; Golden-Marx et al., 2022). Interestingly, the decrease in that results from the inclusion of M14, is reduced from 0.04 dex to 0.02 dex as we extend our measurement to larger radii, which may reflect that M14 is more strongly correlated with the light in the inner portion of the total BCG+ICL profile, as investigated in Section 5.2.
Additionally, although Figure 2 does not present the measurements, the values given in Table 3 are all greater than 13 from 0. Thus, the M14-stellar mass correlation persists out to 300 kpc. However, decreases as we progress to larger radii, which can be seen by the separation between the solid lines shown in Figure 1.
The presence of a non-zero along with the reduction in and increase in allow us to conclude that M14 is a latent parameter when the SMHM relation is measured out to the ICL. For further verification, we randomly reassign the M14 values. If a non-zero value is measured for , this would suggest an intrinsic, non-physical, correlation between these parameters. However, as shown in Table 6, for each bin, the shuffled value is equivalent to 0. Moreover, the values of , , and are in excellent agreement between the versions with the randomly assigned M14 (Table 6) and when M14 is not incorporated (Table 3). Thus, any change in parameter measurements results solely from a correlation between the BCG+ICL stellar mass and M14. Going forward, we include M14 in all measurements discussed in this analysis (though we provide the posteriors when M14 is not accounted for in Tables 3, 4, and 5).
5.2 Comparing the SMHM relation of the BCG to the SMHM relation of outskirts
Although the correlation between the stellar mass and the halo mass increases as the BCG+ICL light profile is extended outward, as discussed in Golden-Marx et al. (2023) and Huang et al. (2022), the strongest correlation between stellar mass and halo mass has been measured when the core of the BCG is excluded. Here, we investigate the inclusion of M14 and use 5 non-overlapping radial apertures, 0-30 kpc, 30 kpc-80 kpc (representative of the ICL transition region (Zhang et al., 2024)), 80 kpc-150 kpc, 150 kpc-225 kpc, 225 kpc-300 kpc. We note that the data set used in this part of the analysis is a slightly different sample as what was used in Section 5.1 because we require that the aperture have a non-zero stellar mass in each of the 5 apertures, so there are some where we can measure a non-zero total light profile (starting from the core), but not an individual aperture mass (i.e., there are some clusters where the data may not be deep enough to allow for measurements of the large outer apertures).
In Figure 3, we plot the the data used to measure the SMHM relation, the stellar mass vs. the halo mass coloured using M14. Although the stellar mass range contained within each aperture varies, we fix the y-axis, to highlight the changing value. In each subplot, we overlay the median values of stellar mass and halo mass for the three previously defined M14 bins. The stellar mass-M14 correlation persists strongly in both the 0-30 kpc and 30 kpc-80 kpc apertures. Beyond this aperture, both the overall distribution and any trend with M14 becomes dominated by the noise, as demonstrated by the solid lines which begin to overlap.
We visually compare how using the core of the BCG impacts the , , and parameters in Figure 4. For , we measure the steepest slope in the radial range 30-80 kpc. This slope that is almost 0.07 (15%) steeper than any version measured when the light from the core of the BCG is used, which supports that the light contained within the ICL is strongly correlated with the cluster’s halo mass. Moreover, this is similar to the result found in Golden-Marx et al. (2023) that the slope is peaked in the radial aperture of 30 kpc-50 kpc and also agrees with results presented in Huang et al. (2022) that find the strongest correlation with halo mass using the radial range of 50 kpc - 100 kpc. Additionally, we note that the light from the outskirts (beyond 150 kpc) is almost as strongly correlated with the dark matter halo mass as the BCG’s core, though with a significantly larger scatter.
The trend measured for is nearly identical to what is found in Golden-Marx et al. (2023), where is significantly lower when the light from the core of the BCG is included. Since BCG cores can be treated as a standard candle (e.g., Lauer et al., 2007), this is unsurprising. However, we remind the reader that our estimate for is an upper limit because we assume the same uncertainty on the stellar mass for each of the radial bins, calibrated for the BCG 0-30 kpc region. If we are underestimating the uncertainty in the stellar mass of the ICL, then we we would be overestimating .
Lastly, the most interesting and surprising result comes from , which is statistically more than 5 from 0 for the 0-30 kpc, 30 kpc-80 kpc, and 80 kpc-150 kpc apertures. This means that the stellar mass-M14 correlation persists out to 150 kpc, though its strength decreases drastically beyond 80 kpc. Moreover, the value of is statistically equivalent in the BCG’s core and the BCG+ICL transition region (30 kpc-80 kpc). Since M14 is independent of the light from this radial regime, this may offer further support that the light in the transition region is also growing as a result of BCG hierarchical assembly. Additionally, we note that at large radii (r > 150 kpc), we detect no correlation between stellar mass and M14, which may lead to the decrease in measured when the BCG’s core is included and shows that for that dataset , the correlation between M14 and stellar mass is coming from the inner region (between 0-80 kpc or 0-150 kpc) of the BCG+ICL system.
6 Physical Meaning of the Correlation between the Magnitude Gap and the ICL
6.1 The SMHM Relation
Figures 1 and 3 illustrate the connection between M14 and the ICL as inferred from the SMHM relation. Table 3 and Figure 4 quantify that the ICL and M14 correlate, such that for certain radial bins, the inclusion of M14 yields a tighter SMHM relation. However, as shown in Figure 4, the strength of this correlation weakens as we extend the ICL out to larger apertures and stops being statistically significant at radii beyond 150 kpc.
Figure 4 and Table 3 also highlight that the strength of the correlation between the BCG+ICL light/stellar mass and M14 is within 1 for the 0-30 kpc, 0-80 kpc, and 30 kpc - 80 kpc apertures. Moreover, both when the core is and is not included, declines at radii beyond 80 kpc. While the parameters are similar, the radial range that features the strongest correlation between the stellar mass and the underlying halo mass is the BCG+ICL transition region (30 kpc - 80 kpc). Although this aperture does not have the lowest , the value ( = 0.187) is in agreement with the values assumed when M14 is not incorporated (). Therefore, from a statistical standpoint, there is a trade-off between the lower when the core is included and the higher when it is not. Additionally, we note that the inclusion of M14 decreases by 0.034 dex and increases by 0.13, in the transition region, which is the largest change in both parameters for one of the unique apertures when incorporating M14 into the SMHM relation.
We remind the reader that the stellar mass measured between 30 kpc-80 kpc and M14 are independent. Thus, this strong parameter informs the BCG+ICL system’s formation. Following the results of Golden-Marx & Miller (2018), M14 can be thought of as a tracer for hierarchical growth, such that clusters with larger M14 values are more evolved and possibly older systems. Thus, the presence of such a strong M14 stratification in the BCG+ICL transition region suggests that the hierarchical growth of the BCG is not confined to the core, but rather extends to the BCG+ICL transition regime. Moreover, it is plausible that a greater fraction of the stellar mass growth from this hierarchical merging is deposited within the transition region, which may lead to the strong correlation. Assuming this scenario, the BCG+ICL transition region then grows predominately through the major/minor mergers of the BCG, while the ICL may grow through tidal stripping of satellite galaxies (as shown by the detected colour gradient (e.g., Montes & Trujillo, 2018; DeMaio et al., 2018; Contini et al., 2019; Golden-Marx et al., 2023)). Moreover, because we use radial apertures to separate the ICL and BCG, it is possible that due to projection effects, the correlation with the ICL is even stronger as some fraction of the ICL is contained inside 30 kpc.
Although redshift evolution is beyond the scope of this analysis, we present an alternative formation scenario. Golden-Marx et al. (2022) find the stellar mass contained within the core of the BCG is constant over this redshift range. Thus, Golden-Marx et al. (2022) posit that the growth must be occurring at larger radial apertures, possibly within the ICL. However, this growth was not detected in Golden-Marx et al. (2023), though that analysis focused on the 50 kpc - 300 kpc aperture (and did not account for M14), which may have washed out any evolution signal due to the increase in noise. Thus, the correlation with M14 may suggest recent growth is not occurring in the far outskirts of the ICL, but rather in the BCG+ICL transition region, as Zhang et al. (2024) find preliminary evidence for.
6.2 The ICL Fraction vs. Magnitude Gap
Along with the correlation between the ICL and M14 highlighted via the SMHM relation, we measure the ICL fraction as a function of M14 since both parameters have been posited to correlate with the dynamical state of the cluster. We measure three unique ICL fractions defined by the following equations.
Equation 7 is the ratio of the ICL light to the total light within the cluster (ICL + satellite galaxies, excluding the BCG) measured between 30 kpc and 150 kpc.
Equation 8 compares the BCG+ICL light to the total light of the cluster measured within 150 kpc.
Equation 9 compares the ICL light to the total light in the cluster within 150 kpc.
For this portion of the analysis, we stack the measurements of the ICL in 5 M14 bins (20% quintiles), and then measure the mean ICL fraction and M14 in each bin. For this analysis, we do not further bin the data by halo mass as Zhang et al. (2024), using the same ICL measurements, detect no correlation between ICL fraction and richness. Moreover, we note that there is no correlation between M14 and halo mass in our measurements.
As shown in Figure 5, we for the first time using observations find that clusters characterized by a large M14 have significantly larger ICL fractions. Each ICL fraction increases by between 50-75% when comparing low M14 (M14 < 1.0) and high M14 (M14 > 2.0) clusters. As a result of hierarchical merging, we expect a fraction of stars will be ejected into the ICL while another fraction will increase the BCG stellar mass. Thus, the M14 and ICL fraction correlation directly results from BCG hierarchical assembly. Therefore, it follows the increase in the fraction as a function of M14 is much larger when the BCG’s light is included (comparing the middle and lower panels), as the BCG’s stellar mass increases due to hierarchical assembly. Moreover, following the interpretation of M14 from Golden-Marx & Miller (2018) and M12 from Vitorelli et al. (2018) – clusters with larger M14 values form earlier and have had more time to evolve – we posit that this may mean that clusters characterized by larger ICL fractions are dynamically older and that as clusters evolve, the BCG+ICL system (and as a result M14 and ICL fraction) grows. Additionally, though not shown, similar to Zhang et al. (2024), we find that the ICL fraction is dependent on aperture and decreases as the ICL is extended out to 300 kpc
Previous work by Jiménez-Teja et al. (2018), Dupke et al. (2022), and Jiménez-Teja et al. (2024) have found that the ICL fraction has an inverse correlation with the dynamical state of the cluster, such that relaxed clusters are characterized by low ICL fractions, while merging clusters are characterized by larger ICL fractions. Moreover, studying fossil group galaxies, Zarattini et al. (2015) suggest that systems characterized by larger M12 values are dynamically relaxed. Additionally, in cosmological hydrodynamic simulations Yoo et al. (2024) find that clusters with larger M12 values are more relaxed. Furthermore, Yoo et al. (2024) also find a strong correlation between M12 and BCG+ICL fraction, similar to what is shown in the middle panel of Figure 5. Although we use M14, as shown in Golden-Marx & Miller (2018), there is no statistical difference between these parameters due to the measurement uncertainty. Thus, the results shown in Figure 5 are in excellent agreement with the simulations from Yoo et al. (2024). Lastly, recent observational work by Casas et al. (2024) also find that applying a high M14 criteria (M14 > 1.25) better identifies dynamically relaxed clusters.
This literature comparison is not a perfect. Jiménez-Teja et al. (2018), Dupke et al. (2022), and Jiménez-Teja et al. (2024) use X-ray measurements, measure the ICL using CICLE (e.g., Jiménez-Teja et al., 2018), and do not estimate the ICL fraction within a fixed aperture, Zarattini et al. (2015) focus on lower halo mass fossil galaxies, Yoo et al. (2024) use simulations the majority of which are galaxy groups, and Casas et al. (2024) also analyze an X-ray sample and do not consider the ICL. However, the excellent agreement with the simulations from Yoo et al. (2024) suggests that both ICL fraction and M14 may correlate with dynamical state such that larger M14 values and higher ICL fractions are typical of dynamically relaxed systems. However, given this contradiction to results from Jiménez-Teja et al. (2018), Dupke et al. (2022), and Jiménez-Teja et al. (2024) we posit two alternatives. First, X-ray relaxation measurements are made on large cluster scale observations. However, M14 depends only on the cluster’s central region. Therefore, it is possible that while M14 traces the hierarchical assembly of the BCG, it is unrelated to the cluster’s overall dynamical state because it is characteristic of just the mergers within the core. Second, Zhang et al. (2024) find that the ICL fraction is dependent on the choice of radial aperture, and thus it is possible that the measurements from Jiménez-Teja et al. (2018), Dupke et al. (2022), and Jiménez-Teja et al. (2024) are not directly comparable to our own. However, further investigation, using simulations and consistent ICL measurement methods and definitions are required to gain greater insight into this apparent discrepancy.
7 Conclusion
In this work, our analysis focuses on characterizing how the ICL and M14 are related using the SMHM relation as both measurements are related to BCG hierarchical assembly. We use DES Year 6 data to characterize the SMHM relation over the redshift range . Throughout the analysis we treat the BCG and ICL as a linked system using radial cuts to define certain regions. We summarize the primary results of this analysis as follows.
When the BCG+ICL stellar mass is measured from the core out to progressively larger apertures, , the correlation between stellar mass and M14 at fixed halo mass is non-zero out to large radii. Using our Bayesian infrastructure, we find that the inclusion of M14 increases by 0.1-0.15 and decreases by 0.01-0.04 dex (depending on the choice of aperture). Thus, M14 is a latent variable in the SMHM relation. However, the impact of M14 decreases at larger radii.
When measuring the SMHM relation in unique, non-overlapping apertures, we find that the strongest correlation between stellar mass and halo mass is in the 30 kpc - 80 kpc bin. This result follows from the “two-phase” formation of BCGs and agrees with results from Huang et al. (2022) and Golden-Marx et al. (2023). Moreover, increases by 30% with the inclusion of M14.
When using unique apertures, the stellar mass - M14 correlation is statistically significant in the 0-30 kpc, 30 kpc - 80 kpc, and 80 kpc - 150 kpc apertures. decreases significantly beyond 80 kpc. Most importantly, is equivalent in the 0-30 kpc, 0-80 kpc, and 30 kpc - 80 kpc apertures. Since M14 contains no information about the light in the 30 kpc - 80 kpc aperture, this correlation highlights that the stellar mass in this region likely grows as a result of BCG hierarchical assembly. Thus, it is plausible that the BCG+ICL transition region grows as a result of mergers, while the outer ICL grows through processes such as tidal stripping.
We for the first time observationally measure a strong correlation between the mean ICL fraction and M14 value such that clusters characterized by a larger M14 (M14 > 2.0) have ICL fractions that are more than 50% higher than clusters with a small M14 (M14 < 1.0). As M14 grows as a result of mergers, this correlation follows from BCG hierarchical assembly.
Going forward, there remain many research questions we aim to explore to improve our observational understanding and characterisation of the BCG+ICL. In particular, understanding how the BCG+ICL stellar mass and radial distribution evolves with redshift. In future analysis using the Euclid Mission (as demonstrated by Kluge et al. (2024)) and the Rubin Observatory’s LSST, we plan to extend this type of analysis out to higher redshifts () to investigate the formation and evolution of the BCG+ICL system and how its correlation with halo mass evolves. Moreover, given the correlation between M14 and ICL fraction, we suggest that characterizing large M14 systems in Euclid and LSST, or targetting them with JWST, may provide information about the earliest appearances of the ICL. Additionally, we aim to determine whether the correlation between M14 and BCG stellar mass persists out to high redshift and whether this is intrinsic (i.e., is it in place prior to ex-situ growth) or results entirely from hierarchical formation.
Acknowledgements
Author Contributions: We would like to acknowledge everyone who made this work possible. JBGM developed the Bayesian infrastructure that was used in this analysis, made the measurements of each parameter, did the analysis of the ICL and M14, and was the principle author responsible for writing the manuscript. YZ developed the infrastructure used to measure the ICL in the DES data, produced the ICL light profiles used in this analysis, provided guidance, and aided in paper writing and discussion. RLCO provided useful insights throughout this work and aided in paper writing and discussion. BY assisted with our use of the DES data and construction of the ICL pipeline. MESP and MH provided detailed and useful feedback as DES internal reviewers. The remaining authors have made contributions to this paper that include, but are not limited to, the construction of DECam and other aspects of collecting the data; data processing and calibration; developing broadly used methods, codes, and simulations; running the pipelines and validation tests; and promoting the science analysis.
JBGM would like to thank Nina Hatch, the members of the NottICL research group, and the participants in the ISSI ICL working group for many useful discussions about the nature of ICL and measurement techniques. JBGM gratefully acknowledges support from the Leverhulme Trust.
Funding: Funding for the DES Projects has been provided by the U.S. Department of Energy, the U.S. National Science Foundation, the Ministry of Science and Education of Spain, the Science and Technology Facilities Council of the United Kingdom, the Higher Education Funding Council for England, the National Center for Supercomputing Applications at the University of Illinois at Urbana-Champaign, the Kavli Institute of Cosmological Physics at the University of Chicago, the Center for Cosmology and Astro-Particle Physics at the Ohio State University, the Mitchell Institute for Fundamental Physics and Astronomy at Texas A&M University, Financiadora de Estudos e Projetos, Fundação Carlos Chagas Filho de Amparo à Pesquisa do Estado do Rio de Janeiro, Conselho Nacional de Desenvolvimento Científico e Tecnológico and the Ministério da Ciência, Tecnologia e Inovação, the Deutsche Forschungsgemeinschaft and the Collaborating Institutions in the Dark Energy Survey.
The Collaborating Institutions are Argonne National Laboratory, the University of California at Santa Cruz, the University of Cambridge, Centro de Investigaciones Energéticas, Medioambientales y Tecnológicas-Madrid, the University of Chicago, University College London, the DES-Brazil Consortium, the University of Edinburgh, the Eidgenössische Technische Hochschule (ETH) Zürich, Fermi National Accelerator Laboratory, the University of Illinois at Urbana-Champaign, the Institut de Ciències de l’Espai (IEEC/CSIC), the Institut de Física d’Altes Energies, Lawrence Berkeley National Laboratory, the Ludwig-Maximilians Universität München and the associated Excellence Cluster Universe, the University of Michigan, NSF’s NOIRLab, the University of Nottingham, The Ohio State University, the University of Pennsylvania, the University of Portsmouth, SLAC National Accelerator Laboratory, Stanford University, the University of Sussex, Texas A&M University, and the OzDES Membership Consortium.
Based in part on observations at Cerro Tololo Inter-American Observatory at NSF’s NOIRLab (NOIRLab Prop. ID 2012B-0001; PI: J. Frieman), which is managed by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the National Science Foundation.
The DES data management system is supported by the National Science Foundation under Grant Numbers AST-1138766 and AST-1536171. The DES participants from Spanish institutions are partially supported by MICINN under grants ESP2017-89838, PGC2018-094773, PGC2018-102021, SEV-2016-0588, SEV-2016-0597, and MDM-2015-0509, some of which include ERDF funds from the European Union. IFAE is partially funded by the CERCA program of the Generalitat de Catalunya. Research leading to these results has received funding from the European Research Council under the European Union’s Seventh Framework Program (FP7/2007-2013) including ERC grant agreements 240672, 291329, and 306478. We acknowledge support from the Brazilian Instituto Nacional de Ciênciae Tecnologia (INCT) do e-Universo (CNPq grant 465376/2014-2).
This manuscript has been authored by Fermi Research Alliance, LLC under Contract No. DE-AC02-07CH11359 with the U.S. Department of Energy, Office of Science, Office of High Energy Physics.
Data Availability
The data underlying this article will be shared on reasonable request to the corresponding author.
Appendix A Verification Tests
In Tables 4 and 5, we present the posterior distributions when we divide our sample into two subsets based on the redMaPPer central galaxy probability (). We use two samples with (Table 4) and (Table 5). When comparing these posterior distributions to those presented in Table 3, we see that all measurements are consistent within 1. This provides additional evidence that our results are not biased by miscentring or incorrect BCG identification, as BCGs with low are more likely to be miscentred or misidentified as the BCG.
| Data | Inner Radius | Outer Radius | ||||
| Core | 0 | 30 | – | |||
| Core | 0 | 80 | – | |||
| Core | 0 | 150 | – | |||
| Core | 0 | 225 | – | |||
| Core | 0 | 300 | – | |||
| Core | 0 | 30 | ||||
| Core | 0 | 80 | ||||
| Core | 0 | 150 | ||||
| Core | 0 | 225 | ||||
| Core | 0 | 300 | ||||
| No Core | 0 | 30 | – | |||
| No Core | 30 | 80 | – | |||
| No Core | 80 | 150 | – | |||
| No Core | 150 | 225 | – | |||
| No Core | 225 | 300 | – | |||
| No Core | 0 | 30 | ||||
| No Core | 30 | 80 | ||||
| No Core | 80 | 150 | ||||
| No Core | 150 | 225 | ||||
| No Core | 225 | 300 |
| Data | Inner Radius | Outer Radius | ||||
| Core | 0 | 30 | – | |||
| Core | 0 | 80 | – | |||
| Core | 0 | 150 | – | |||
| Core | 0 | 225 | – | |||
| Core | 0 | 300 | – | |||
| Core | 0 | 30 | ||||
| Core | 0 | 80 | ||||
| Core | 0 | 150 | ||||
| Core | 0 | 225 | ||||
| Core | 0 | 300 | ||||
| No Core | 0 | 30 | – | |||
| No Core | 30 | 80 | – | |||
| No Core | 80 | 150 | – | |||
| No Core | 150 | 225 | – | |||
| No Core | 225 | 300 | – | |||
| No Core | 0 | 30 | ||||
| No Core | 30 | 80 | ||||
| No Core | 80 | 150 | ||||
| No Core | 150 | 225 | ||||
| No Core | 225 | 300 |
In Table 6, we present the posterior distributions when M14 is randomly assigned, removing any existing correlation between M14 and aperture stellar mass. For each version, we find that is equivalent to 0 and that the measured posteriors are in excellent agreement with the values provided in Table 3 when M14 is not accounted for. The discrepancy between the shuffled results (Table 6) and those of the real data (Table 3) along with the nested nature of our MCMC model further highlight that M14 is indeed a latent parameter, the correlation between aperture stellar mass and M14 is real, and including M14 yields a better fit to the underlying observational data.
| Data | Inner Radius | Outer Radius | ||||
| Core | 0 | 30 | ||||
| Core | 0 | 80 | ||||
| Core | 0 | 150 | ||||
| Core | 0 | 225 | ||||
| Core | 0 | 300 | ||||
| No Core | 0 | 30 | ||||
| No Core | 30 | 80 | ||||
| No Core | 80 | 150 | ||||
| No Core | 150 | 225 | ||||
| No Core | 225 | 300 |